A roler has some $4/lb coffee and some $8/lb coffee that she will mix to make 50 lb of $5.60/lb coffee. How many pounds of each should be used?
Answer by stanbon(75887)

A roler has some $4/lb coffee and some $8/lb coffee that she will mix to make 50 lb of $5.60/lb coffee. How many pounds of each should be used?
Let amount of $4 coffee be "x" lbs.
Then amount of $8 coffee is "50-x" lbs.
Value of this $4 coffee is 4x dollars.
Value of the $8 coffee is 8(50-x)=400-8x dollars
Value of the mix is 5.6(50)= 280 dollars
EQUATION:
value + value = value
4x+400-8x=280
-4x=-120
x=30 lbs. (amount of $4 coffee)
50-x=20 lbs. (amount of $8 coffee)
Cheers,
Stan H.